### v2

# Manual2Skill: Learning to Read Manuals and Acquire Robotic Skills for Furniture Assembly Using Vision-Language Models

Abstract
Humans possess an extraordinary ability to understand and execute complex manipulation tasks by interpreting abstract instruction manuals. For robots, however, this capability remains a substantial challenge, as they cannot interpret abstract instructions and translate them into executable actions. In this paper, we present Manual2Skill, a novel framework that enables robots to perform complex assembly tasks guided by high-level manual instructions. Our approach leverages a Vision-Language Model (VLM) to extract structured information from instructional images and then uses this information to construct hierarchical assembly graphs. These graphs represent parts, subassemblies, and the relationships between them. To facilitate task execution, a pose estimation model predicts the relative 6D poses of components at each assembly step. At the same time, a motion planning module generates actionable sequences for real-world robotic implementation. We demonstrate the effectiveness of Manual2Skill by successfully assembling several real-world IKEA furniture items. This application highlights its ability to manage long-horizon manipulation tasks with both efficiency and precision, significantly enhancing the practicality of robot learning from instruction manuals. This work marks a step forward in advancing robotic systems capable of understanding and executing complex manipulation tasks in a manner akin to human capabilities. Project Page: https://owensun2004.github.io/Furniture-Assembly-Web/
Figure 1: Overview of Manual2Skill Framework. We propose Manual2Skill, which learns manipulation skills from manuals, enabling robots to understand and execute complex manipulation tasks in a manner akin to humans. The green region showcases the input of our pipeline: the pictures of the assembly manual and real parts. The blue region depicts our pipeline: 1) a Vision-Language Model (VLM) generates a Hierarchical Assembly Graph, 2) a per-step pose estimation module predicts the 6D-poses of components, and 3) a motion planning and execution module controls the robot arms to assemble the furniture autonomously.
11footnotetext: Equal contribution.
I Introduction
Humans can learn manipulation skills from instructions in images or texts; for example, people can assemble IKEA furniture or LEGO models by following a manual’s instructions. This ability enables humans to efficiently acquire long-horizon manipulation skills from sketched instructions. In contrast, robots typically learn such skills through imitation learning [59] or reinforcement learning [43], both of which require significantly more data and computation. Replicating the human ability to transfer abstract manuals to real-world actions remains a significant challenge for robots. Manuals are typically designed for human understanding, using simple schematic diagrams and symbols to convey manipulation processes. This abstraction makes it difficult for robots to comprehend such instructions and derive actionable manipulation strategies  [32, 49, 48]. Developing a method for robots to effectively utilize human-designed manuals would greatly expand their capacity to tackle complex, long-horizon tasks while reducing the demand of collecting extensive demonstration data.
Manuals inherently encode the structural information of complex tasks. They decompose high-level goals into mid-level subgoals and capture task flow and dependencies, such as sequential steps or parallelizable subtasks. For example, furniture assembly manuals guide the preparation and combination of components and ensure that all steps follow the correct order  [32]. Extracting this structure is crucial for robots to replicate human-like understanding and manage complex tasks effectively [19, 33]. After decomposing the task, robots need to infer the specific information for each step, such as the involved components and their spatial relationships. For example, in cooking tasks, the instruction images and texts may involve selecting ingredients, tools, and utensils and arranging them in a specific order  [38]. Finally, robots need to generate a sequence of actions to complete the task, such as grasping, placing, and connecting components. Previous works have tried to leverage sketched pictures [42] or trajectories [15] to learn manipulation skills but are always limited to relatively simple tabletop tasks.
In this paper, we propose Manual2Skill, a novel robot learning framework that is capable of learning manipulation skills from visual instruction manuals. This framework can be applied to automatically assemble IKEA furniture, a challenging and practical task that requires complex manipulation skills. As illustrated in Figure 1, given a set of manual images and the real furniture parts, we first leverage a vision language model to understand the manual and extract the assembly structure, represented as a hierarchical graph. Then, we train a model to estimate the assembly poses of all involved components in each step. Finally, a motion planning module generates action sequences to move selected components to target poses and executes them on robots to assemble the furniture.
In summary, our main contributions are as follows:
  • We propose Manual2Skill, a novel framework that leverages VLM to learn robotic skills from manuals, enabling a generalizable assembly pipeline for IKEA furniture.
  • We introduce a hierarchical graph generation pipeline that utilizes VLM to extract structured information for assembly tasks. Our pipeline facilitates real-world assembly and extends to other assembly applications.
  • We define a novel assembly pose estimation task within the learning-from-manual framework. We predict the 6D poses of all involved components at each assembly step to meet real-world assembly requirements.
  • We evaluate our method on four real items of IKEA furniture, demonstrating its effectiveness and applicability in real-world assembly tasks.
II Related Work
II-A Furniture Assembly
Part assembly is a long-standing challenge with extensive research exploring how to construct a complete shape from individual components or parts [6, 13, 20, 27, 29, 36, 53, 46, 45]. Broadly, we can categorize part assembly into geometric assembly and semantic assembly. Geometric assembly relies solely on geometric cues, such as surface shapes or edge features, to determine how parts mate together [6, 53, 37, 10]. In contrast, semantic assembly primarily leverages high-level semantic information about the parts to guide assembly process [13, 20, 27, 29, 45].
Furniture assembly is a representative semantic assembly task, where each part has a predefined semantic role (e.g., a chair leg or a tabletop), and the assembly process follows intuitive, common-sense relationships (e.g., a chair leg must be attached to the chair seat). Previous studies on furniture assembly have tackled different aspects of the problem, including the motion planning [41], multi-robot collaboration [25], and assembly pose estimation [29, 58, 30]. Researchers have developed several datasets and simulation environments to facilitate research in this domain. For example, Wang et al. [49], Liu et al. [32] introduced IKEA furniture assembly datasets containing 3D models of furniture and structured assembly procedures derived from instruction manuals. Additionally, Lee et al. [27] and Yu et al. [58] developed simulation environments for IKEA furniture assembly, while Heo et al. [16] provides a reproducible benchmark for real-world furniture assembly. However, existing works typically focus on specific subproblems rather than addressing the entire assembly pipeline. In this work, we aim to develop a comprehensive framework that learns the sequential process of furniture assembly from manuals and deploys it in real-world experiments.
II-B VLM Guided Robot Learning
Vision Language Models (VLMs) [57] have been widely used in robotics to understand the environment [17] and interact with humans [39]. Recent advancements highlight VLMs’ potential to enhance robot learning by integrating vision and language information, enabling robots to perform complex tasks with greater adaptability and efficiency [18]. A potential direction is the development of the Vision Language Action Model (VLA Model) that can generate actions based on the vision and language inputs [2, 23, 3, 44]. However, training such models requires vast amounts of data, and they struggle with long-horizon or complex manipulation tasks. Another direction is to leverage VLMs to guide robot learning by providing high-level instructions and perceptual understanding. VLMs can assist with task descriptions [17, 18], environment comprehension [19], task planning [47, 56, 62], and even direct robot control [28]. Additionally, Goldberg et al. [14] demonstrates how VLMs can assist in designing robot assembly tasks. Building on these insights, we explore how VLMs can interpret abstract manuals and extract structured information to guide robotic skill learning for long-horizon manipulation tasks.
II-C Learning from Demonstrations
Learning from demonstration (LfD) has achieved promising results in acquiring robot manipulation skills [12, 64, 7]. For a broader review of LfD in robotic assembly, we refer to  Zhu and Hu [65]. The key idea is to learn a policy that imitates the expert’s behavior. However, previous learning methods often require fine-grained demonstrations, like robot trajectories [7] or videos [22, 40, 21]. Collecting these demonstrations is often labor-intensive and may not always be feasible. Some works propose to learn from coarse-grained demonstrations, like the hand-drawn sketches of desired scenes [42] or rough trajectory sketches [15]. These approaches reduce dependence on expert demonstrations and improve the practicality of LfD. However, they are mostly limited to tabletop manipulation tasks and do not generalize well to more complex, long-horizon assembly problems. In this work, we aim to extend LfD beyond these constraints by tackling a more challenging assembly task using abstract instruction manuals.
III Problem Formulation
Given a complete set of 3D assembly parts and its assembly manual, our goal is to generate a physically feasible sequence of robotic assembly actions for autonomous furniture assembly. Manuals typically use schematic diagrams and symbols designed to depict step-by-step instructions in an abstract format that is universally understandable. We define the manual pages as a set of $N$ images. =$\mathcal{I}=\{I_{1},I_{2},\cdots,I_{N}\}$, where each image $I_{i}$ illustrates a specific step in the assembly process, such as the merging of certain parts or subassemblies
The furniture consists of $M$ individual parts $\mathcal{P}=\{P_{1},P_{2},\cdots,P_{M}\}$. A part is an individual element in $\mathcal{P}$ that remains disconnected from other parts until assembly. A subassembly is any partially or fully assembled structure that forms a proper subset of $\mathcal{P}$ (for example, $\{P_{1},P_{2}\}$). The term component encompasses both parts and subassemblies.
Given the manual and 3D parts, the system generates an assembly plan. Each step corresponds to a manual image and specifies the involved parts and sub-assemblies, their spatial 6D poses, and the assembly actions or motion trajectories required for execution.
IV Technical Approach
Our approach automates furniture assembly by leveraging the VLM to interpret IKEA-style manuals and guide robotic execution. Given a visual manual and physical parts in a pre-assembly scene, a VLM generates a hierarchical assembly graph, defining which parts and subassemblies are involved in each step. Next, a per-step pose estimation model predicts 6D poses for each component using a manual image and the point clouds of involved components. Finally, for assembly execution, the estimated poses are transformed into the robot’s world frame, and a motion planner generates a collision-free trajectory for part mating.
This paper shows an overview of our framework in Fig. 2. We describe the VLM-guided assembly hierarchical graph generation in Section IV-A, followed by per-step assembly pose estimation in Section IV-B and assembly action generation based on component relationships in Section IV-C.
Figure 2: Framework Overview. (1) GPT-4o [1] is queried with manual pages to generate a sequential assembly plan, represented as a hierarchical assembly graph. (2) The furniture components’ point clouds and corresponding manual images are processed by a pose estimation module to predict target poses for each component. (3) The system sequentially executes the assembly by planning and performing robotic actions based on the hierarchical assembly graph and estimated poses.
IV-A VLM Guided Hierarchical Assembly Graph Generation
This section demonstrates how VLMs can interpret IKEA-styled manuals to generate high-level assembly plans. Given a manual and a real-world image of furniture parts (pre-assembly scene image), a VLM predicts a hierarchical assembly graph. We show one example in Fig. 2. In this graph, leaf nodes represent atomic parts, while non-leaf nodes denote subassemblies. We structure the graph in multiple layers, where each layer contains nodes representing parts or subassemblies involved in a single assembly step (corresponding to one manual image). The directed edges from the children to a parent node indicate that the system assembles the parent node from all its children nodes. Additionally, we add edges between equivalent parts, denoting these parts are identical($e.g.$ four legs of a chair). Representing the assembly process as a hierarchical graph can decomposes the assembly into sequential steps while specifying necessary parts and subassemblies. We give the formal definition of the hierarchical graph in Section -J. We achieve this in two stages: Associating Manuals with Real Parts and Identifying Parts needed in Each Image.
IV-A1 VLM Capabilities and General Prompt Structure
The task is inherently complex due to the diverse nature of input images. Manuals are typically abstract sketches, whereas pre-assembly scene images are high-resolution real-world images. Such diversity requires advanced visual recognition and spatial reasoning across varied image domains, which are strengths of VLMs due to their training on extensive, internet-scale datasets. We demonstrate the effectiveness of VLMs for this task in Section V-A and Section -D.
Every VLM prompt consists of two components:
  • Image Set: This includes all manual pages and the real-world pre-assembly scene image. Unlike traditional VLM applications in robotics [23, 18], which process a single image, our method requires multi-image reasoning.
  • Text Instructions: These instructions provide a task-specific context, guiding the model in interpreting the image set. The instructions range from simple directives to Chain-of-Thought reasoning [51]. All instructions incorporate in-context learning examples, specifying the required output format—be it JSON, Python code, or natural language. This structure is essential to our multi-stage pipeline, ensuring well-structured, interpretable outputs that seamlessly integrate into subsequent stages.
IV-A2 Stage I: Associating Real Parts with Manuals
Given the manual’s cover sketch of the assembled furniture and the pre-assembly scene image, the VLM aims to associate physical parts with the manual. The VLM achieves this by predicting the roles of each physical part through semantically interpreting the manual’s illustrations. This process involves analyzing spatial, contextual, and functional cues in the manual illustrations to enable a comprehensive understanding of each physical part. Our method follows CoT [51] and Least-to-Most [63] prompting for better accuracy.
To enhance part identification, we employ Set of Marks [55] and GroundingDINO [31] to automatically label parts on the pre-assembly scene image with numerical indices. The labeled scene image and manual sketch form the Image Set. Text instructions consist of a brief context explanation for the association task of predicting the roles of each physical part, accompanied by in-context examples of the output structure:
{name, label, role}
For example, in Figure 2 In Stage I Output, we describe the chair’s seat as name: seat frame, label: [2], role: for people sitting on a chair, the seat offers essential support and comfort and is positioned centrally within the chair’s frame. Here, ‘[2]’ indicates that this triplet corresponds to the physical part labeled with index 2 in the pre-assembly scene image. This triplet format enhances interpretability and ensures consistency by structuring all outputs into the same data format. We use the Image Set and Text Instructions as the input prompt for the VLM (specifically GPT-4o [1]) and query it once to generate real assignments for all physical parts. We then use these labels as leaf nodes in the hierarchical assembly graph.
We can obtain equivalent parts through these triplets. When two physical parts share the same geometric shapes, their triplets only differ by label. For example, in Figure 2 Stage I Output, {name: side frame, label: [0], role:…} and {name: side frame, label: [1], role:…}—these two parts are considered equivalent. Understanding equivalent part relationships is crucial for downstream modules, as demonstrated by our ablation experiments(see Section -C).
IV-A3 Stage II: Identify Involved Parts in Each Step
This stage focuses on identifying the particular parts and subassemblies involved in each manual page. The VLM achieves this by reasoning through the illustrated assembly steps, using the triplets and the labeled pre-assembly scene from the previous stage as supporting hints.
In practice, we observe that irrelevant elements in the manual (e.g., nails, human figures) can distract the VLM. We use the cropped manual images from [49], where only the furniture parts and subassemblies are reserved to focus the VLM’s attention (Figure 2 Stage II Image Set), significantly improving performance (see Section -E for details).
The manual pages, combined with the labeled pre-assembly scene from the previous stage, form the Image Set. The Text Instructions use a Chain-of-Thought prompt to guide the VLM in identifying parts and subassemblies step by step and includes in-context examples that clarify the structured output format: a pair consisting of (Step N, Labeled Parts Involved). The bottom left output of Figure 2 provides an example of this format. Together, the Image Set and Text Instructions compose the input prompt for GPT-4o, which generates pairs for all assembly steps using a single query.
As shown in Fig. 2, the system outputs nested lists. We then transform these lists, along with the equivalent parts, into a hierarchical graph. Using this assembly graph, we traverse all non-leaf nodes and explore various assembly orders. Formally, a feasible assembly order is an ordered set of non-leaf nodes, ensuring that a parent node appears only after all its child nodes. A key advantage of the hierarchical graph representation is its flexibility—since the assembly sequence is not unique, it allows for parallel assembly or strategic sequencing.
IV-B Per-step Assembly Pose Estimation
Given an assembly order, we train a model to estimate the poses of components (parts or subassemblies) at each step of the assembly process. At each step, the model inputs the manual image and the point clouds of the involved components, predicting their target poses to ensure proper alignment. To support this task, we construct a dataset for sequential pose estimation. For a detailed description, see Section -A.
Given each component’s point cloud (obtained from real-world scans or our dataset), we first center it by translating its centroid to the origin. Next, we apply Principal Component Analysis (PCA) to identify the dominant object axes, which define a canonical coordinate frame. The most dominant axes serve as the reference frame, ensuring a shape-driven and consistent orientation that remains independent of arbitrary coordinate systems.
The dataset we create provides manual images, point clouds, and target poses for each component in the camera frame of the corresponding manual image(following [29]). For an assembly step depicted in the manual image $\mathcal{I}_{i}$, the inputs to our model include: (1) the manual image $\mathcal{I}_{i}$; (2) the point clouds of all involved components. The output is the target pose $T\in SE(3)$ for each component represented in the camera frame of $I_{i}$.
IV-B1 Model Architecture
Note that the number of components at each step is not fixed, depending on the subassembly division of the furniture. Our pose estimation model consists of four parts: an image encoder $\mathcal{E}_{I}$, a point cloud encoder $\mathcal{E}_{P}$, a cross-modality fusion module $\mathcal{E}_{G}$, and a pose regressor $\mathcal{R}$.
We first feed the manual image $I$ into the image encoder to get an image feature map $\mathbf{F}_{I}$.
$\mathbf{F}_{I}=\mathcal{E}_{I}(I)$ (1)
Then, we feed the point clouds into the point cloud encoder to get the point cloud feature for each component.
$\{\mathbf{F}_{j}\}=\mathcal{E}_{P}(\{P\}_{j})$ (2)
In order to fuse the multi-modality information from the manual image and the point cloud features, we leverage a GNN [54] to update the information for each component. We consider the manual image feature and component-wise point cloud features as nodes in a complete graph, employing a GNN to update the information for each node.
$\mathbf{F}_{I}^{\prime},\{\mathbf{F}_{j}^{\prime}\}=\mathcal{E}_{G}(\mathbf{F}_{I},\{\mathbf{F}_{j}\})$ (3)
where $\mathbf{F}_{I}^{\prime},\{\mathbf{F}_{j}^{\prime}\}$ are updated image and point cloud features.
Finally, we feed the updated point cloud features as input into the pose regressor to get the target pose for each component.
$T_{j}=\mathcal{R}(\mathbf{F}_{j}^{\prime})$ (4)
IV-B2 Loss Function
t jointly considers pose prediction accuracy and point cloud alignment, following [60, 30]. The first term penalizes errors in the predicted SE(3) transformation, while the second measures the distance between predicted and ground truth point clouds. To account for interchangeable components, we compute the loss across all possible permutations of equivalent parts and select the minimum loss as the final training objective. We provide further details on the loss formulation and training strategy in  Section -B.
IV-C Robot Assembly Action Generation
IV-C1 Align Predicted Poses with the World Frame
At each assembly step, the previous stage predicts each component’s pose in the camera frame of the manual image. However, real-world robotic systems operate in their world frame, requiring a 6D transformation between these coordinates. Consider two components, A and B. The predicted target poses in the camera frame are denoted as ${}^{I_{i}}\hat{\mathcal{T}}_{a}$ and ${}^{I_{i}}\hat{\mathcal{T}}_{b}$. Meanwhile, our system can collect the current 6D pose of part A in the world frame, represented as ${}^{W}\mathcal{T}_{a}$. To align ${}^{I_{i}}\hat{\mathcal{T}}_{a}$ to ${}^{W}\mathcal{T}_{a}$, we compute the 6D transformation matrix ${}^{W}_{I_{i}}\mathcal{T}$, which maps the camera frame to the world frame.
${}^{W}\mathcal{T}_{a}={}^{W}_{I_{i}}\mathcal{T}{}^{I_{i}}\hat{\mathcal{T}}_{a}$ (5)
Using the same transformation ${}^{W}_{I_{i}}\mathcal{T}$, we compute the assembled target pose of part B (and all remaining components) in the world frame.
${}^{W}\mathcal{T}_{b}={}^{W}_{I_{i}}\mathcal{T}{}^{I_{i}}\hat{\mathcal{T}}_{b}$ (6)
This transformation accurately maps predicted poses from the manual image frame to the robot’s world frame, ensuring precise assembly execution.
IV-C2 Assembly Execution
Once our system determines the target poses of each component in the world frame for the current assembly step, it grasps each component and generates the required action sequences for assembly.
Part Grasping
After scanning each real-world part, we obtain the corresponding 3D meshes for each part. We employ FoundationPose [52], and the Segment Anything Model (SAM) [24] to obtain the initial poses of all parts in the scene.
Given the pose and shape of each part, we design heuristic grasping methods tailored to the geometry of individual components. While general grasping algorithms such as GraspNet [11] are viable, grasping is beyond the scope of this work. Instead, we employ heuristic grasping strategies specifically designed for structured components in assembly tasks. For stick-shaped components, we grasp the centroid of the object after identifying its longest axis for stability. For flat and thin-shaped components, we use fixtures or staging platforms to securely position the object, allowing the robot to grasp along the thin boundary for improved stability. We provide further details on these grasping methods in Section -G.
Part Assembly Trajectory
Once the robot arm grasps a component, it finds a feasible, collision-free path to predefined robot poses (anchor poses). At these poses, the 6D pose of the grasped component is recalculated in the world frame, leveraging the FoundationPose  [52] and the Segment Anything Model (SAM)[24]. The system then plans a collision-free trajectory to the component’s target pose. We use RRT-Connect [26] as our motion planning algorithm. All collision objects in the scene are represented as point clouds and fed into the planner. Once the planner finds a collision-free path, the robot moves along the planned trajectory.
Assembly Insertion Policy
Once the robot arm moves a component near its target pose, the assembly insertion process begins. Assembly insertions are contact-rich tasks that require multi-modal sensing (e.g., force sensors and closed-loop control) to ensure precise alignment and secure connections. However, developing closed-loop assembly insertion skills is beyond the scope of this work and will be addressed in future research. In our current approach, human experts manually perform the insertion action.
V Experiments
In this section, we perform a series of experiments aimed at addressing the following questions.
  • Q1: Can our proposed hierarchical assembly graph generation module effectively extract structured information from manuals? (see Section V-A)
  • Q2: Can the per-step pose estimation be applicable to different categories of furniture and outperform previous settings? (see Section V-B)
  • Q3: How effective is the proposed framework in the assembly of furniture with manual guidance? (see Section V-C)
  • Q4: Can this pipeline be applied to real-world scenarios?(see Section V-D)
  • Q5: Can this pipeline be extended to other assembly tasks? (see Section V-E)
  • Q6: How should we determine and evaluate the key design choices of each module? (ablation experiments, see Sections -E and -C)
In addition, we have included a comprehensive set of prompts utilized in the VLM-guided hierarchical graph generation process in Section -K
V-A Hierarchical Assembly Graph Generation
In this section, we evaluate the performance of our VLM-guided hierarchical assembly graph generation approach. Specifically, we assess Stage II: Identifying Parts in Each Image using the IKEA-Manuals dataset [49]. We provide the rationale for excluding Stage I evaluation in Section -H.
TABLE I: Success Rate Across Task Complexity ($\uparrow$)
Number of Parts 2$\sim$4 5$\sim$6 7$\sim$8 9+ Average
SingleStep 35.7 16.7 0 0 10.8
GeoCluster 21.4 2.8 0 0 3.9
Ours (Scene Variations) 78.6 50.0 12.5 0 31.4
Ours 78.6 55.6 8.3 0 32.4
Furniture Count 14 36 24 28
Figure 3: Qualitative results. Our method significantly outperforms the baselines. SingleStep fails on moderately complex furniture, while GeoCluster generates physically impossible subassemblies (highlighted in red). In contrast, our approach closely aligns with the ground truth.
Figure 4: Pre-Assembly Scene Variations. (Left) original pre-assembly scene. (Middle) parts randomly shuffled along the ground plane. (Right) parts randomly rotated in-place.
Experiment Setup. The IKEA-Manuals dataset [49] includes 102 furniture items, each with IKEA manuals, 3D parts, and assembly plans represented as trees in nested lists. For each item, we load its 3D parts into Blender and render two images: one depicting the original pre-assembly scene with neatly arranged parts, and another showing a scene variation where parts are arbitrarily perturbed (e.g., rotated and shuffled) along a ground plane (see Figure 4). This randomization introduces diversity, better simulating real-world pre-assembly scenarios where parts may be disorganized.
Each rendered image of the pre-assembly scene, along with the manual, is processed by the VLM through the stages outlined in Section IV-A to generate a hierarchical assembly graph. Since we represent our graph as a nested list, we align our notation with the assembly tree notation used in IKEA-Manuals [49]. In this subsection, we refer to our generated assembly graph as the predicted tree.
Evaluation Metric. We use Success Rate criterion, which measures the proportion of the predicted tree that exactly matches the ground-truth tree. We consider a predicted tree exactly matched if all its nodes contain the same set of children nodes as their ground-truth counterparts. All equivalent parts are seen as identical when computing the tree matching.
Baselines. We compare our VLM-based method against two heuristic approaches introduced in IKEA-Manuals [49].
  • SingleStep predicts a flat, one-level tree with a single parent node and $n$ leaf nodes.
  • GeoCluster employs a pre-trained DGCNN [50] to iteratively group furniture parts with similar geometric features into a single assembly step. Compared to SingleStep, it generates deeper trees with more parent nodes and multiple hierarchical levels.
Results. As shown in LABEL:tab:assembly_plan, understanding and interpreting manuals is a challenging task. Our VLM‑guided approach effectively handles furniture manuals with up to 6 parts, a significant breakthrough where baselines struggle even in simpler cases. This $\leq 6$ parts threshold reflects current VLM capabilities in complex visual-spatial reasoning tasks. Our framework is designed to be scalable, benefiting from rapid VLM advancements. We anticipate performance improvements as more powerful VLMs emerge. LABEL:tab:assembly_plan also highlights the VLM’s generalization ability: pre-assembly scene variations in Figure 4 minimally impact assembly graph generation performance, with average success rates drop within 1% across all 102 furniture items. Figure 3 provides qualitative results for two furniture items, illustrating the advantages of our approach in greater detail.
For ablation studies on the necessity of using and segmenting manuals, see Section -E. Failure cases are further explained in Section -F. Additional results and prompt templates are in Sections -D and -K, respectively.
V-B Per-step Assembly Pose Estimation
TABLE II: Quantitative Results of Pose Estimation.
[…]
Method Chair Lamp Table Chair Lamp Table Chair Lamp Table Chair Lamp Table
Li et al. [29] 1.847 1.865 1.894 0.247 0.278 0.318 0.243 0.396 0.519 0.268 0.121 0.055
Mean-Max Pool 0.434 1.118 1.059 0.087 0.187 0.200 0.046 0.229 0.280 0.457 0.199 0.107
Ours 0.202 0.826 0.953 0.042 0.153 0.172 0.027 0.189 0.276 0.868 0.240 0.184
Data Preparation. We select three categories of furniture items from PartNet [34]: chair, table, and lamp. For each category, we select 100 furniture items and generate 10 parts selection and subassembly division for each piece of furniture. To generate the assembly manual images, we render diagrammatic images of parts at 20 random camera poses using Blender’s Freestyle functionality. We provide more details about it in Section -A. In general, we generate 12,000 training and 5,200 testing data pieces for each category.
Training Details. For the Image Encoder $\mathcal{E}_{I}$, we selected the encoder component of DeepLabV3+, which includes MobileNet V2 as the backbone and the atrous spatial pyramid pooling (ASPP) module. We made this choice because DeepLabV3+ leverages atrous convolutions on the basis of Auto Encoder, enabling the model to capture multi-scale structures and spatial information effectively [4, 5]. It generates a multi-channel feature map from the image $I$, and we use mean-max pool [61] to derive a global vector $\mathbf{F}_{I}\in\mathbb{R}^{256}$ from the feature map. For the Point Clouds Encoder $\mathcal{E}_{P}$, we use the encoder part of PointNet++ [35]. For each part and subassembly, we extract a part-wise feature $\mathbf{F}_{j}\in\mathbb{R}^{256}$. For the GNN $\mathcal{E}_{G}$, we use a three-layer graph transformer [8]. The pose regressor $\mathcal{R}$ is a three-layer MLP. We provide more details of the mean-max pool for the image feature and our training hyperparameter setting in Section -B.
Baselines. We evaluate the performance of our method on our proposed per-step assembly pose estimation dataset. We compare our method with two baselines:
  • Li et al. [29] proposed a pipeline for single image guided 3D object pose estimation.
  • Mean-Max Pool is a variant of our method, replacing GNN with a mean-max pool trick, similar to our approach of obtaining a one-dimensional vector from a multi-channel feature map, with details in Section -B.
Evaluation Metrics. We adopt comprehensive evaluation metrics to assess the performance of our method and baselines.
  • Geodesic Distance (GD), which measures the shortest path distance on the unit sphere between the predicted and ground-truth rotations.
  • Root Mean Squared Error (RMSE), which measures the Euclidean distance between the predicted and ground-truth poses.
  • Chamfer Distance (CD), which calculates the holistic distance between the predicted and the ground-truth point clouds.
  • Part Accuracy (PA), which computes the Chamfer Distance between the predicted and the ground truth point clouds; if the distance is smaller than 0.01m, we count this part as “correctly placed”.
Results. As shown in Table II, our method outperforms  Li et al. [29] and the mean-max pool variant in all evaluation metrics and on three furniture categories. We attribute this to the effectiveness of our multi-modal feature fusion and GNN in capturing the spatial relationships between parts. We also provide qualitative results for each furniture category in Figure 5.
Ablation. To assess the impact of equivalent parts, guided image, and per-step data about subassemblies, we perform ablation studies on these components. We present the details and results in Section -C.
Figure 5: Qualitative results on three furniture categories. We observe better pose predictions than baselines.
Figure 6: Qualitative Evaluation on real IKEA furniture items. This figure illustrates the assembly process of various IKEA furniture items, including FLISAT, VARIERA, SUNDVIK, and KNAGGLIG, with our approach. For each item, we display the manual images, per-step 3D parts pose estimation results, and real-world assembly outcomes.
V-C Overall Performance Evaluation
We evaluate the overall performance of our method by assembling furniture models in a simulation environment. We implement the evaluation process in the PyBullet [9] simulation environment and test the entire pipeline. We source all test furniture models from the IKEA-Manuals dataset [49]. Given these manuals along with 3D parts, we generate the pre-assembly scene images as described in  IV-C, and our pipeline generates the hierarchical graphs. Then, we traverse the hierarchical graph to determine the assembly order. Following this sequence and the predicted 6D poses of each component, we implement RRT-Connect [26] in simulation to plan feasible motion paths for the 3D parts and subassemblies, ensuring they move towards their target poses. Note that, in this experiment, we focus on object-centric motion planning and omit robotic execution in our framework.
Baselines. As the first to propose a comprehensive pipeline for furniture assembly, there is no direct baseline for comparison. So we design a baseline method that uses previous work  [29] to estimate the poses of all parts, with the guidance of an image of the fully assembled furniture, and adopt a heuristic order to assemble all parts. Specifically, given the predicted poses of all parts, we can calculate the distance between each pair of parts. The heuristic order is defined as follows: starting from a random part, we find the nearest part to it and assemble it, then successively find the nearest part to the assembled parts until we assemble all parts.
Evaluation Metrics. We adopt the assembly success rate as the evaluation metric and define the following situations as a failure: 1) A part is placed at a pose that is too far from the ground truth pose. 2) A part collides with other parts when moving to the estimated pose. In other words, the RRT-Connect algorithm [26] finds no feasible path when mating it with other parts. 3) We place a part that is not near any other components, causing it to suspend in midair after each assembly step.
TABLE III: Success Rate on 4 Furniture Categories()($\uparrow$)( ↑ )
Method Bench Chair Table Misc Average
Li et al. [29]+Heuristic 0.00 0.39 0.11 0.00 0.30
Ours 0.67 0.61 0.44 0.50 0.58
Results. We evaluate the overall performance on 50 furniture items from the IKEA-Manual dataset [49], each consisting of fewer than seven parts. These items fall into four categories (Bench, Chair, Table, Misc), and we report the success rate for each in Table III.
Our system successfully assembles 29 out of 50 furniture pieces, whereas the baseline method assembles only 15. Our framework achieves a success rate of 58%, demonstrating the effectiveness of our proposed framework. The most common failure occurs when the VLM fails to generate a fully accurate assembly graph, leading to misalignment between the point cloud and the instruction manual images used for pose estimation.
Figure 7: Real-World Setup. We use two UFactory xArm6 for assembly and a RealSense D435 camera for pose estimation.
V-D Real-world Assembly Experiments
To evaluate the feasibility and performance of our pipeline, we conducted experiments in the real world using four IKEA furniture items: Flisat (Wooden Stool), Variera (Iron Shelf), Sundvik (Chair), and Knagglig (Box). Figure 7 illustrates our real-world experiment setup. We show the manual images, per-step pose estimation results, and real-world assembly process in Figure 6. We also attach videos of the real-world assembly process in the supplementary material. For detailed implementation of our real-world experiments, please check Section -G. We evaluated all the assembly tasks with target poses provided by three different methods: Ground truth Pose, Mean-Max Pool (see Section V-B), and our proposed approach. The Ground truth Pose method uses the ground truth poses for each part to assemble the furniture. We use the Average Completion Rate (ACR) as the evaluation criterion and calculate it as follows:
$ACR=\frac{1}{N}\sum_{j=1}^{N}\frac{S_{j}}{S_{\text{total}}}$ (7)
where $N$ is the total number of trials, $S_{j}$ is the number of steps completed in trial $j$, and $S_{\text{total}}$ denotes the total number of steps in the task.
We perform each task over 10 trials with varying initial 3D part poses. We present the results in Table IV, showing that our method outperforms the baseline and achieves a high success rate in real-world assembly tasks.
These findings underscore the practicality and effectiveness of our approach for real-world implementation. The primary failure mode arises from planning limitations, particularly in handling complex obstacles. Failures occur when the RRT-Connect algorithm cannot find a feasible trajectory when the planned path results in collisions with the robotic arm or surrounding objects or due to suboptimal grasping poses. To improve robustness in real-world scenarios, we plan to develop a low-level policy for adaptive motion refinements—a topic we leave for future work.
TABLE IV: Real World Success Rate ($\uparrow$) over 10 trials.
Method FLISAT VARIERA SUNDVIK KNAGGLIG
Oracle Pose 72.5 85.0 80.0 90.0
Mean-Max Pool 52.5 61.7 40.0 70.0
Ours 60.0 80.0 68.0 85.0
V-E Generalization to Other Assembly Tasks
Figure 8: Pipeline Extension Beyond Furniture Assembly.
We design Manual2Skill as a generalizable framework capable of handling diverse assembly tasks with manual instructions. To assess its versatility, we evaluate the VLM-guided hierarchical graph generation method across three distinct assembly tasks, each varying in complexity and application domain. These include: (1) Assembling a Toy Car Axle (a low-complexity task with standardized components, representing consumer product assembly), (2) Assembling an Aircraft Model (a medium-complexity task, representing consumer product assembly), and (3) Assembling a Robotic Arm (a high-complexity task involving non-standardized components, representing research & prototyping assembly).
For the toy car axle and aircraft model, we sourced 3D parts from [46] and reconstructed pre-assembly scene images using Blender. We manually crafted the manuals in their signature style, with each page depicting a single assembly step through abstract illustrations. For the robotic arm assembly, we used the Zortrax robotic arm [66], which includes pre-existing 3D parts and a structured manual. These inputs were then processed through the VLM-guided hierarchical graph generation pipeline (described in Sec. V-A), yielding assembly graphs as shown in Figure 8. This zero-shot generalization achieves a success rate of $100\%$ over five trials per task. The generated graphs align with ground-truth assembly sequences, confirming the generalization of our VLM-guided hierarchical graph generation across diverse manual-based assembly tasks and highlighting its potential for broader applications.
VI Limitations
This paper explores the acquisition of complex manipulation skills from manuals and introduces a method for automated IKEA furniture assembly. Despite this progress, several limitations remain. First, our approach mainly identifies the objects that need assembly but overlooks other details on the manual, such as grasping position markings and precise connector locations (e.g., screws). Integrating a vision-language model (VLM) module to extract this information could significantly enhance robotic insertion capabilities. Second, the method does not cover the automated execution of fastening mechanisms, like screwing or insertion actions, which depend heavily on force and tactile sensing signals. We leave these challenges as directions for future work.
VII Conclusion
In this paper, we address the issue of learning complex manipulation skills from manuals, which is essential for robots to execute such tasks based on human-designed instructions. We propose Manual2Skill, a novel framework that leverages VLM to understand manuals and learn robotic manipulation skills from manuals. We design a pipeline for assembling IKEA furniture and validate its effectiveness in real scenarios. We also demonstrate that our method extends beyond the task of furniture assembly. This work represents a significant step toward enabling robots to learn complex manipulation skills with human-like understanding. It could potentially unlock new avenues for robots to acquire diverse complex manipulation skills from human instructions.